OpenStudy (anonymous):

write 3ab^-2c^-1 with positive exponents

OpenStudy (kropot72):

\[10x ^{-2}=\frac{10}{x ^{2}}\] You can make use of this rule.

OpenStudy (anonymous):

i don't get it

OpenStudy (anonymous):

\[\sqrt[3]{xy^2}\]

OpenStudy (anonymous):

i am totally baffles by the one above too,,,

OpenStudy (kropot72):

To eliminate the negative exponent of x I used the following process: \[10x ^{-2}=\frac{10x ^{-2}}{1}\] Next multiply the numerator an the denominator of the fraction by x^2 \[\frac{10x ^{-2}\times x ^{2}}{x ^{2}}=\frac{10x ^{0}}{x ^{2}}=\frac{10}{x ^{2}}\]

OpenStudy (anonymous):

where did u get the 10 x to start with???

OpenStudy (kropot72):

The choice of 10x^-2 was just an example of the process to eliminate a negative exponent (without actually giving the answer to your question).

OpenStudy (kropot72):

Good!

OpenStudy (anonymous):

3ab/1 -2 oh sisgh

OpenStudy (anonymous):

what do u do with the exta exponents?

OpenStudy (anonymous):

is it just 3 over all the variables?

OpenStudy (kropot72):

\[3ab ^{-2}c ^{-1}=\frac{3a}{b ^{2}c}\]

OpenStudy (anonymous):

oh becaause the a doesn't have an exponet,,